Ally is 8 years older than Dan. Dan is half Ken's age. Put them in order from oldest to youngest
Answer by greenestamps(13203)



There is not enough information.

We know that both Ally and Ken are older than Dan. (Ally is 8 years older than Dan; Dan is half Ken's age.)
But depending on the exact ages, Ken could be either older of younger than Ally.

Dan 2; Ally 2+8=10; Ken 2*2=4 --> oldest to youngest = Ally, Ken, Dan
Dan 12; Ally = 12+8=20; Ken 12*2=24 --> oldest to youngest = Ken, Ally, Dan